# EnvGen: Generating and Adapting Environments via LLMs for Training Embodied Agents

### v1

@@ Abstract
Recent state-of-the-art approaches for embodied learning via interaction directly employ large language models (LLMs) as agents to determine the next steps in an environment. Due to their world knowledge and reasoning capabilities, LLM agents achieve stronger performance than previous smaller agents based on reinforcement learning (RL); however, frequently calling LLMs is slow and expensive. This begs an interesting question: Instead of directly employing LLMs as embodied agents, can we use LLMs’ reasoning capabilities to adaptively create training environments to help smaller embodied RL agents learn useful skills that they are weak at? In this work, we propose EnvGen, a novel framework to address this question. First, we prompt an LLM to generate training environments that allow agents to quickly learn different tasks in parallel. Concretely, the LLM is given the task description and environment simulator objectives that the agents should learn and is then asked to generate a set of environment configurations (e.g., different terrains, items initially given to agents, chance of finding certain objects, etc.). Next, we train a small RL agent in a mixture of the original and LLM-generated environments. Then, we enable the LLM to continuously adapt the generated environments to progressively improve the skills that the agent is weak at, by providing feedback to the LLM in the form of the agent’s performance. We demonstrate the usefulness of EnvGen with comprehensive experiments in Crafter and Heist game environments. We find that a small RL agent trained with EnvGen can outperform SOTA methods, including a GPT-4 agent, and learns long-horizon tasks significantly faster. We also show qualitatively how the LLM adapts training environments to help improve RL agents’ weaker skills over time. Additionally, EnvGen is substantially more efficient as it only uses a small number of LLM calls (e.g., 4 in total), whereas LLM agents require one or more LLM calls per step (resulting in thousands of LLM calls per episode). Lastly, we present detailed ablation studies for EnvGen’s design choices.
@@ 1 Introduction
To address this question, we propose EnvGen, a novel framework where an LLM adaptively generates training environments to teach smaller embodied RL agents. We aim to generate environments that can create various conditions (e.g., have different terrains or some subgoals are already achieved) so that agents can learn different skills in parallel and obtain more frequent rewards for challenging long-horizon tasks than in the original environment. As shown in Fig. 1 (c), EnvGen iterates over multiple training cycles, each consisting of the following four steps.
@@
Step 2: We train the RL agent in multiple LLM-generated environments (i.e., LLM environments), so that it can learn different useful skills in parallel.
Step 3: We first train the RL agent in the original environment to mitigate overfitting to the LLM environments. Then we measure the current RL agent’s performance in different tasks in the original environment to check which skills/tasks the agent is still weak at.
@@
We study the usefulness of EnvGen in different game environments: Crafter (Hafner, 2022) and Heist (Cobbe et al., 2020). In the Crafter environment, a simple PPO-based (Schulman et al., 2017) lightweight (M parameters) RL agent trained with our LLM-generated environments outperforms strong baselines including a GPT-4 based agent that queries an LLM multiple times at every step, and RL agents that use extensive pretraining (e.g., 150M steps vs. less than 1M steps for us). When compared to just training longer in the original Crafter environment, an RL agent trained with EnvGen achieves significant improvements on the overall score and long-horizon tasks. In Heist, we also show that our LLM-generated environments can improve overall agent performance and training stability. We also show a qualitative study on how the LLM adapts training environments to help improve RL agents’ weaker skills over time. Finally, we provide comprehensive analysis and ablation studies of the design choices of EnvGen, including EnvGen vs. longer training in the original environment, adaptively updating LLM-generated environments vs. a fixed environment, different LLMs for generating environments, frequency of environment updates, the number of LLM-generated environments, and the mixture ratio between the original and LLM environment during training.
@@ 2 EnvGen: Generating and Adapting Environments via LLMs for Training Embodied Agents
We propose EnvGen, a novel framework where an LLM adaptively generates training environments to train smaller embodied RL agents, enabling them to accomplish various tasks within an environment, particularly long-horizon tasks. During the training process, the LLM is given feedback (in the form of the agent’s performance) and can adaptively update the training environments to progressively focus on improving the tasks that the agent is weak at. In the following, we first explain why it is challenging to explore long-horizon tasks in open-world games (Sec. 2.1). Then we explain our method details, including how we generate environments and how agents are trained in the generated and original environments (Sec. 2.2).
@@ 2.2 EnvGen Method Details
Step 1: Generate training environments with an LLM.
As illustrated in step 1 of Fig. 2, we use an LLM (e.g., GPT-4 (OpenAI, 2023a)111We use GPT-4-1106-Preview (i.e., GPT-4 Turbo).) to first generate custom training environment configurations222We find that N=4 works well; see Table 7 for details. that can cover various objectives and skills that are required in the original environment. The following describes the LLM input prompt components used to create environment configurations.
@@
Game/simulator details: We provide a list of objectives that need to be achieved in the environment (e.g., “collect coal, collect iron, etc.” for Crafter); a list of which simulator settings can be controlled (e.g., terrain, agent inventory); and a list of constraints/rules that the simulator has (e.g., “skeletons only spawn in mountains; cows only spawn in grass; …” for Crafter).
@@
The obtained environment configurations are then rendered in the game’s simulator. Fig. 2 presents the summary of input prompt and output environments from the GPT-4 model. We provide more prompt details in the appendix.
Step 2: Train a small RL agent in the LLM-generated environments.
As shown in step 2 of Fig. 2, we first train the small RL agent in the LLM-generated environments. Concretely, we train the agent in the LLM environments for total steps in parallel.
Step 3: Train and measure the RL agent’s performance in the original environment.
It is important to note that the goal of EnvGen is to improve the RL agent’s performance in the original environment, instead of the performance only in the LLM environments. To help the RL agent effectively adapt to the original environment and provide the LLM with the current agent’s performance as feedback, we train the agent and measure its performance in the original environment, as shown in step 3 of Fig. 2. First, to mitigate the overfitting to LLM environments, we train the agent in the original environment for steps.333We find that works well; see Table 8 for details. Next, to find the skills that the RL agent needs to improve at, we test the agent in the original environment, without any parameter updates. Concretely, we measure individual success rates for each environment task (e.g., Crafter achievements). The agent performance is summarized (in step 4) and is provided to LLM as feedback (in step 1) to adapt training environments in the next cycle. Moreover, importantly, to obtain a more calibrated estimation of agent performance, we calculate the average and variance of the task-specific scores by testing agents with multiple random seeds (i.e., 12). Note this performance measuring takes minimal time and computation compared to the actual training steps (e.g., 30x faster in our experiments) as it does not involve backpropagation.
Step 4: Send feedback to LLM to adapt environments (to focus on weak skills).
We provide the LLM with the agent’s performance from the original environment (measured in step 3), as feedback for updating LLM environments. Concretely, we list the agent’s average task-specific success rate in percentages along with one standard deviation (e.g., “ collect coal: 38% 6%, defeat skeleton: 10% 4% ”), as shown in step 4 of Fig. 2. In step 1 of the next cycle, the LLM can adaptively generate new environments (by using the agent’s performance as feedback) to better help the RL agent learn the skills it is weak at (e.g., defeat skeleton). EnvGen iterates this four-step training cycle times.
@@ 3.1 Evaluated Benchmarks and Training Details
Crafter.
Crafter (Hafner, 2022) is an open-world 2D survival game focused on evaluating a broad range of agent capabilities (see Fig. 3). Crafter features 22 achievements that an agent can unlock during an episode of play. Some achievements can be unlocked in a few steps (e.g., collect wood, collect sapling, etc.), but other achievements, such as make iron pickaxe or collect diamond, require many training/exploration steps and several prerequisite achievements to be unlocked (see Fig. 3 b). For example, to make a stone pickaxe, an agent must first collect enough wood to make a table and a wooden pickaxe, then go collect stone and return to the table (or collect more wood to make a new one) and then construct the stone pickaxe. As an agent progresses, these prerequisite achievements become increasingly compounded (e.g., an iron pickaxe requires a table + a stone pickaxe + a furnace + some coal + some wood), making it incredibly difficult to complete all achievements. The score for Crafter is computed as the geometric mean of individual success rates of each achievement for each episode it is completed within 1M training steps: , where is the average success rate of the th achievement across all episodes that occurred during training.
For EnvGen setup, we use training cycles during agent training (see Table 6 for ablation of having a different number of cycles). Each cycle uses 0.12M LLM-generated environment steps (i.e., Crafter steps, see step 2 in Fig. 2) and 0.12M Crafter steps ( step 3 in Fig. 2) and then we train for 1M steps in Crafter. In total, we train for 1.96M steps ((0.12M + 0.12M) 4 + 1M). Note that in order to maintain a fair score comparison to baselines, we do not count any achievement during our training cycle for score calculation since the training scores derived from LLM environments and the original environment are not directly comparable. Instead, we only take into account the achievements from the last 1M training steps in Crafter for the score calculation. We also experiment with giving the baseline model additional original environment steps to match the number of EnvGen steps (i.e., an additional 0.96M steps) to ensure that EnvGen is not better simply because of more steps. We report the average performance with 30 runs (= 3 different initial LLM-generated Crafter environments 10 different random seeds).
Heist.
@@ 3.2 Agent Architectures
Our base RL agent.
For both Crafter and Heist, we test the EnvGen framework with a simple (CNN + linear layer) and lightweight (5M) agent used in Moon et al. (2023), which is slightly modified from the agent architecture used in IMPALA (Espeholt et al., 2018). Following Moon et al. (2023), we train the agent with a PPO (Schulman et al., 2017) objective. At every step, the agent takes an RGB image (surroundings for Crafter, entire maze for Heist) as input and outputs the value estimates and policy (action probability). See Fig. 3 (a) and Fig. 4 (a) for example visual inputs for agents. We provide additional model implementation details in the appendix.
Baseline methods.
For Crafter, we compare our method to two groups of recent baselines – (1) methods that use frequent (i.e., more than thousands of) LLM calls during training or inference: SPRING (Wu et al., 2023) (based on GPT-4 (OpenAI, 2023a)) and ELLM (Du et al., 2023) (based on Codex (Chen et al., 2021)) and (2) methods that do not use an LLM during training or inference: DreamerV3 (Hafner et al., 2023), MuZero + SPR (Walker et al., 2023), LSTM-SPCNN (Stanić et al., 2023), PPO (Schulman et al., 2017), and Achievement Distillation (AD) (Moon et al., 2023). For Heist, we compare against the PPO agent. For the PPO and AD agents, we follow the implementation of Moon et al. (2023). See appendix for the PPO/AD agent details.
@@ 4 Results and Analysis
We demonstrate the usefulness of the EnvGen method with comprehensive experiments and analysis. We first compare RL agents trained with EnvGen to different baseline methods on Crafter, an open-world game with 22 hierarchical achievements (Sec. 4.1). Next, we present a detailed analysis of the improvements that training with EnvGen environments can give RL agents on long-horizon tasks (Sec. 4.2). Then, we analyze how the LLM-based environment adaptation can help an RL agent progressively improve the skills that the agent is weak at (Sec. 4.3). Moreover, we also compare an RL agent’s performance on Heist (a maze navigation game), with and without EnvGen environments (Sec. 4.4). Lastly, we present various ablation studies on EnvGen design choices (Sec. 4.5).
@@
| LSTM-SPCNN (Stanić et al., 2023) | 135M | 11.7 0.8 | 9.3 0.2 | ||
| DreamerV3 (Hafner et al., 2023) | 201M | 14.8 1.4 | 10.9 0.5 | ||
| MuZero + SPR (Walker et al., 2023) | 150M step PT in Crafter w/ RND reward | 54M | 16.4 1.5 | 12.7 0.4 | |
| SPRING* (Wu et al., 2023) | 9 queries to call GPT-4 per step | 2.7K | Unknown | 27.3 1.2 | 12.3 0.7 |
@@ 4.1 Comparison with State-of-the-art Methods on Crafter Environment
Small RL agent trained with EnvGen outperforms state-of-the-art baselines.
On the Crafter environment (described in Sec. 3.1), we compare a small PPO agent trained in Crafter (i.e., Crafter environments generated with EnvGen) to state-of-the-art baseline methods. As shown in Table 1, we find that a small (4M parameters) PPO agent with EnvGen achieves an average score of 32.2% and significantly outperforms the baselines (and also in terms of the average reward). Note that some baseline agents have many more parameters or pretraining steps such as SPRING (27.3%) that directly employs GPT-4 as agent, and MuZero + SPR (16.4%) that uses 150M pretraining steps. Also, note that our method only uses orders of magnitude fewer LLM calls (only 4) than works like SPRING (2.7K on average) and ELLM (5M). Due to the large number of LLM calls, SPRING costs around $270 USD to run an agent in a single episode, whereas EnvGen only costs a couple of cents total for any number of episodes (see appendix for cost details). Moreover, we find that EnvGen can help train another small RL agent – Achievement Distillation (AD) (Moon et al., 2023) – to achieve an even higher score (35.3%).
@@ 4.2 Detailed Achievement Analysis on Crafter Environment
Next, we analyze where EnvGen improves the overall score by checking individual achievement success rates in detail. For this, we compare the same PPO agent architecture trained with different setups: (1) an agent trained on Crafter for 1.96M steps and (2) an agent trained on Crafter for 0.96M steps (0.24M steps 4 training cycles, see Sec. 2.2) and then trained on Crafter for 1M steps. We measure the success rate (Fig. 5) of each achievement and unlocking speed (Fig. 6) of iron tools in the last 1M training steps, and discuss the results below.
EnvGen helps RL agents to tackle challenging long-horizon achievements.
Fig. 5 shows that training in Crafter improves scores of several achievements. Notably, training in Crafter significantly improves the scores of long-horizon achievements (with many prerequisites; see Fig. 3) such as ‘make stone pickaxe’, ‘make iron pickaxe’, and ‘make iron sword’. Fig. 6 shows that after unlocking the stone pickaxe, the RL agent trained in Crafter is significantly faster in unlocking iron tools compared to the RL agent trained only in Crafter – 40K vs. 135K steps for ‘make iron pickaxe’ and 192K vs. 925K steps for ‘make iron sword’. In the appendix, we additionally show the comparison between two AD (Moon et al., 2023) agents, where the agent trained in Crafter can also improve the success rate of another challenging long-horizon achievement – ‘collect diamond’.

### v2

@@ Abstract
Recent state-of-the-art approaches for embodied learning via interaction directly employ large language models (LLMs) as agents to determine the next steps in an environment. Due to their world knowledge and reasoning capabilities, LLM agents achieve stronger performance than previous smaller agents based on reinforcement learning (RL); however, frequently calling LLMs is slow and expensive. This begs an interesting question: Instead of directly employing LLMs as embodied agents, can we use LLMs’ reasoning capabilities to adaptively create training environments to help smaller embodied RL agents learn useful skills that they are weak at? In this work, we propose EnvGen, a novel framework to address this question. First, we prompt an LLM to generate training environments that allow agents to quickly learn different tasks in parallel. Concretely, the LLM is given the task description and environment simulator objectives that the agents should learn and is then asked to generate a set of environment configurations (e.g., different terrains, items initially given to agents, chances of finding certain objects, etc.). Next, we train a small RL agent in a mixture of the original and LLM-generated environments. Then, we enable the LLM to continuously adapt the generated environments to progressively improve the skills that the agent is weak at, by providing feedback to the LLM in the form of the agent’s performance. We demonstrate the usefulness of EnvGen with comprehensive experiments in Crafter and Heist game environments. We find that a small RL agent trained with EnvGen can outperform SOTA methods, including a GPT-4 agent, and learns long-horizon tasks significantly faster. We also show that using an LLM to adapt environments dynamically outperforms curriculum learning approaches and how the LLM adapts training environments to help improve RL agents’ weaker skills over time. Additionally, EnvGen is substantially more efficient as it only uses a small number of LLM calls (e.g., 4 in total), whereas LLM agents require one or more LLM calls per step (resulting in thousands of LLM calls per episode). We also present detailed analyses of EnvGen’s design choices.
@@ 1 Introduction
To address this question, we propose EnvGen, a novel framework where an LLM adaptively generates training environments to teach smaller embodied RL agents. We aim to generate environments that can create various conditions (e.g., have different terrains or some subgoals are already achieved) so that agents can learn different skills in parallel and obtain more frequent rewards for challenging long-horizon tasks than in the original environment. As shown in Fig. 1 (c), EnvGen iterates over multiple training cycles, each consisting of the following four steps:
@@
Step 2: We first train the RL agent in multiple LLM-generated environments (i.e., LLM environments), so that it can learn different useful skills in parallel.
Step 3: We then train the RL agent in the original environment to mitigate overfitting to the LLM environments. Afterwards, we measure the current RL agent’s performance in different tasks in the original environment to check which skills/tasks the agent is still weak at.
@@
We study the usefulness of EnvGen in different game environments: Crafter (Hafner, 2022) and Heist (Cobbe et al., 2020). In the Crafter environment, a simple PPO-based (Schulman et al., 2017) lightweight (M parameters) RL agent trained with our LLM-generated environments outperforms strong baselines including a GPT-4 based agent that queries an LLM multiple times at every step, and RL agents that use extensive pretraining (e.g., 150M steps vs. less than 1M steps for us). When compared to just training longer in the original Crafter environment and curriculum learning approaches such as easy-to-hard and adversarial environments, an RL agent trained with EnvGen achieves significant improvements on the overall score and long-horizon tasks. In Heist, we also show that our LLM-generated environments can improve overall agent performance and training stability. We also show a qualitative study on how the LLM adapts training environments to help improve RL agents’ weaker skills over time. Finally, we provide comprehensive analysis and ablation studies of the design choices of EnvGen, including dynamically updating LLM environments (i.e., using adaptive environments) vs. curriculum learning methods, frequency of environment updates, EnvGen vs. longer training in the original environment, different LLMs for generating environments, the number of LLM-generated environments, and the mixture ratio between the original and LLM environment during training.
@@ 2 EnvGen: Generating and Adapting Environments via LLMs for Training Embodied Agents
We propose EnvGen, a novel framework where an LLM adaptively generates training environments to train smaller embodied RL agents, enabling them to accomplish various tasks within an environment, particularly long-horizon tasks. During the training process, the LLM is given feedback (in the form of the agent’s performance) and can adaptively update the training environments to progressively focus on improving the tasks that the agent is weak at. In the following, we first explain why it is challenging to explore long-horizon tasks in open-world games (Sec. 2.1). Then we explain our method details, including how we generate environments and how agents are trained in EnvGen (Sec. 2.2).
@@ 2.2 EnvGen Method Details
Step 1: Generate training environments with an LLM. As illustrated in step 1 of Fig. 2, we use an LLM (e.g., GPT-4 (OpenAI, 2023a)) to first generate custom training environment configurations111We find that N=4 works well; see Table 6 for details. that can cover various objectives and skills that are required in the original environment. The following describes the LLM input prompt components used to create environment configurations.
@@
Game/simulator details: We provide a list of objectives that need to be achieved in the environment (e.g., “collect coal, collect iron, etc.” for Crafter); a list of which simulator settings can be controlled (e.g., terrain, agent inventory); and a list of constraints/rules that the simulator has (e.g., “skeletons only spawn in mountains; …” for Crafter).
@@
The obtained environment configurations are then rendered in the game’s simulator. Fig. 2 presents the summary of input prompt and output environments from the GPT-4 model. We provide more prompt details in Appendix F.
Step 2: Train a small RL agent in the LLM-generated environments. As shown in step 2 of Fig. 2, we train the small RL agent in the LLM-generated environments. Concretely, we train the agent in the LLM environments for total steps in parallel.
Step 3: Train and measure the RL agent’s performance in the original environment. It is important to note that the goal of EnvGen is to improve the RL agent’s performance in the original environment, instead of the performance only in the LLM environments. To help the RL agent effectively adapt to the original environment and provide the LLM with the current agent’s performance as feedback, we train the agent and measure its performance in the original environment, as shown in step 3 of Fig. 2. First, to mitigate the overfitting to LLM environments, we train the agent in the original environment for steps.222We find that works well; see Table 7 for details. Next, to find the skills that the RL agent needs to improve at, we test the agent in the original environment, without any parameter updates. Concretely, we measure individual success rates for each environment task (e.g., Crafter achievements). The agent performance is summarized (in step 4) and is provided to LLM as feedback (in step 1) to adapt training environments in the next cycle. Moreover, importantly, to obtain a more calibrated estimation of agent performance, we calculate the average and variance of the task-specific scores by testing agents with multiple random seeds (i.e., 12).
Step 4: Send feedback to LLM to adapt environments (to focus on weak skills). We provide the LLM with the agent’s performance from the original environment (measured in step 3), as feedback for updating LLM environments. Concretely, we list the agent’s average task-specific success rate in percentages along with one standard deviation (e.g., “ collect coal: 38% 6%, defeat skeleton: 10% 4% ”), as shown in step 4 of Fig. 2. In step 1 of the next cycle, the LLM can adaptively generate new environments (by using the agent’s performance as feedback) to better help the RL agent learn the skills it is weak at (e.g., defeat skeleton). EnvGen iterates this four-step training cycle times.
@@ 3.1 Evaluated Benchmarks and Training Details
Crafter. Crafter (Hafner, 2022) is an open-world 2D survival game focused on evaluating a broad range of agent capabilities (see Fig. 3). Crafter features 22 achievements that an agent can unlock during an episode of play. Some achievements can be unlocked in a few steps (e.g., collect wood, collect sapling, etc.), but other achievements, such as make iron pickaxe or collect diamond, require many training/exploration steps and several prerequisite achievements to be unlocked (see Fig. 3 b). For example, to make an iron pickaxe, an agent must first collect enough wood to make a table and a wooden pickaxe, then go collect stone and return to the table (or collect more wood to make a new one) and then construct a stone pickaxe. Then the agent still needs to make a furnace, collect coal, and collect iron before the option to make the iron pickaxe is possible.
For EnvGen setup, we use training cycles during agent training (see Table 3 for ablation of having a different number of cycles). Each cycle uses 0.12M LLM-generated environment steps (i.e., Crafter steps, see step 2 in Fig. 2) and 0.12M Crafter steps ( step 3 in Fig. 2) and then we train for 1M steps in Crafter. In total, we train for 1.96M steps ((0.12M + 0.12M) 4 + 1M). Note that in order to maintain a fair score comparison to baselines, we do not count any achievement during our training cycle for score calculation since the training scores derived from LLM environments and the original environment are not directly comparable. Instead, we only take into account the achievements from the last 1M training steps in Crafter for the score calculation. We also experiment with giving the baseline model additional original environment steps to match the number of EnvGen steps (i.e., an additional 0.96M steps) to ensure that EnvGen is not better simply because of more steps. The score for Crafter is computed as the geometric mean of individual success rates of each achievement for each episode it is completed within 1M training steps: , where is the average success rate of the th achievement across all episodes that occurred during training. We report the average performance with 30 runs (= 3 different initial LLM-generated Crafter environments 10 different random seeds).
@@ 3.2 Agent Architectures
Our base RL agent. For both Crafter and Heist, we test the EnvGen framework with a simple (CNN + linear layer) and lightweight (5M) agent used in Moon et al. (2023), which is slightly modified from the agent architecture used in IMPALA (Espeholt et al., 2018). Following Moon et al. (2023), we train the agent with a PPO (Schulman et al., 2017) objective. At every step, the agent takes an RGB image (surroundings for Crafter, entire maze for Heist) as input and outputs the value estimates and policy (action probability). See Fig. 3 (a) for an agent visual input example. We provide additional model details in Appendix E.
Baseline methods. For Crafter, we compare our method to two groups of recent baselines – (1) methods that use frequent (i.e., more than thousands of) LLM calls during training or inference: SPRING (Wu et al., 2023) (based on GPT-4) and ELLM (Du et al., 2023) (based on Codex (Chen et al., 2021)) and (2) methods that do not use an LLM: DreamerV3 (Hafner et al., 2023), MuZero + SPR (Walker et al., 2023), LSTM-SPCNN (Stanić et al., 2023), PPO (Schulman et al., 2017), and Achievement Distillation (AD) (Moon et al., 2023). For Heist, we compare against the PPO agent. For the PPO and AD agents, we follow the implementation of Moon et al. (2023). See Appendix E for the PPO/AD agent details.
@@ 4 Results and Analysis
We demonstrate the usefulness of the EnvGen method with comprehensive experiments and analysis. We first compare RL agents trained with EnvGen to different baseline methods on Crafter, an open-world game with 22 hierarchical achievements (Sec. 4.1). Next, we present a detailed analysis of the improvements that training with EnvGen environments can give RL agents on long-horizon tasks (Sec. 4.2). Then, we analyze how the LLM-based environment adaptation can help an RL agent progressively improve the skills that the agent is weak at (Sec. 4.3). Lastly, we present various additional analysis including experiments on Heist (a maze navigation game) and ablation studies on EnvGen design choices (Sec. 4.4 and also in the Appendix C).
@@
| LSTM-SPCNN∗ (Stanić et al., 2023) | 135M | 11.7 0.8 | 9.3 0.2 | ||
| DreamerV3∗ (Hafner et al., 2023) | 201M | 14.8 1.4 | 10.9 0.5 | ||
| MuZero + SPR∗ (Walker et al., 2023) | 150M step PT in Crafter w/ RND reward | 54M | 16.4 1.5 | 12.7 0.4 | |
| SPRING* (Wu et al., 2023) | 9 queries to call GPT-4 per step | 2.7K† | Unknown | 27.3 1.2 | 12.3 0.7 |
@@ 4.1 Comparison with State-of-the-art Methods on Crafter Environment
Small RL agent trained with EnvGen outperforms state-of-the-art baselines. On the Crafter environment (described in Sec. 3.1), we compare a small PPO agent trained in Crafter (i.e., Crafter environments generated with EnvGen) to state-of-the-art baseline methods. As shown in Table 1, we find that a small (4M parameters) PPO agent with EnvGen achieves an average score of 32.2% and significantly outperforms the baselines (and also in terms of the average reward). Note that some baseline agents have many more parameters or pretraining steps such as SPRING (GPT-4 agent; 27.3%), and MuZero + SPR (150M pretraining steps; 16.4%). Our method also only uses orders of magnitude fewer LLM calls (only 4) than works like SPRING (2.7K on average) and ELLM (5M), making it much cheaper/more efficient. EnvGen can also work with other RL agents such as Achievement Distillation (AD) (Moon et al., 2023) to achieve an even higher score (35.3%).
@@ 4.2 Detailed Achievement Analysis on Crafter Environment
Next, we analyze where EnvGen improves the overall score by checking individual achievement success rates in detail. For this, we compare the same PPO agent architecture trained with different setups: (1) an agent trained on Crafter for 1.96M steps and (2) an agent trained on Crafter for 0.96M steps (0.24M steps 4 training cycles, see Sec. 2.2) and then trained on Crafter for 1M steps. We measure the success rate (Fig. 4) of each achievement and unlocking speed (Fig. 5) of iron tools in the last 1M training steps.
EnvGen helps RL agents to tackle challenging long-horizon achievements. Fig. 4 shows that training in Crafter improves scores of several achievements. Notably, training in Crafter significantly improves the scores of long-horizon achievements (with many prerequisites; see Fig. 3) such as stone and iron tools. Fig. 5 shows that after unlocking the stone pickaxe, the RL agent trained in Crafter is significantly faster in unlocking iron tools. In Sec. C.1, we also compare two AD agents, and show that Crafter improves the success rate of the most challenging achievement – ‘collect diamond’.